OpenStudy (anonymous):

derivative of f(x)=x^(-1/2)*e^(-x/2) ?

OpenStudy (anonymous):

I know the result but I don't know how we get there: -(e^(-x/2)(x+1)/2x^(3/2)

OpenStudy (anonymous):

\[f(x) = x ^{\frac{ -1 }{ 2 }}\times e ^{\frac{ -x }{ 2 }}\]

OpenStudy (anonymous):

is that your function

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Could you show me the steps please?

OpenStudy (anonymous):

@primeralph He's really good at explaining it

OpenStudy (primeralph):